OpenStudy (lexber):

can someone help me with 3 questions please!!!

OpenStudy (lexber):

simplify \[\sqrt[5]{7}+\sqrt[12]{6}-\sqrt[10]{7}-\sqrt[5]{}\]

OpenStudy (anonymous):

Yea wats ur question but make it quick

OpenStudy (lexber):

@Kitty.10

OpenStudy (lexber):

simplify \[\sqrt{10}\times \sqrt{8}\]

OpenStudy (anonymous):

By is the 5 and all the other numbers exponents can you take a screen shot it makes it much easier for me

OpenStudy (anonymous):

For the second one find the sqrt of 10 and 8 then multiply

OpenStudy (lexber):

\[\sqrt[7]{3}-\sqrt[4]{6}+\sqrt{48}-\sqrt{54}\]

OpenStudy (anonymous):

Again can you take a screen shot of the last one because it can be interpreted in different ways

OpenStudy (anonymous):

Alright for one and three multiple the number INFROUNT by the square root then add or subtract multiple or divide as needed

OpenStudy (anonymous):

Multiply *

OpenStudy (anonymous):

If there isn't a number infrount then just find the square root and add or subtract

OpenStudy (lexber):

wait do i need to do that for each one? i really dont know how to do this so. @Kitty.10

OpenStudy (anonymous):

Yes you do sorry for the late response my phone battery ran out

OpenStudy (anonymous):

@lexber

OpenStudy (anonymous):

So for the first one the square root of seven is 2.6457513 what will it be when you multiply it by five

OpenStudy (anonymous):

@lexber